OpenStudy (anonymous):

Consider the following limit lim 3x/√[(x^2)+3] x→−∞ a) Use the definition of limits at infinity to find values of N that correspond to ε = 0.5 N = b) Use the definition of limits at infinity to find values of N that correspond to ε = 0.1 N = I already found x1 = (ε-3)√[(3)/6ε-ε^2]

OpenStudy (nincompoop):

are you using cauchy?

OpenStudy (anonymous):

no. what's that?

OpenStudy (nincompoop):

nvm

OpenStudy (anonymous):

ok. can you help me with this question?

OpenStudy (anonymous):

I know I put 0.5 and 0.1 in for ε but I when I try to solve it from there I get the wrong answer.

OpenStudy (anonymous):

i dont know what the e is, but to take that limit as it goes to infinity you can divide everything by \[\sqrt{x ^{2}}\] so you will get 3/√[1+3/x] as it goes to infinity and it gives you 3

OpenStudy (anonymous):

\[x _{1}= (\epsilon-3)\frac{ \sqrt{3} }{ 6 \epsilon - \epsilon ^{2}}\]

OpenStudy (anonymous):

the e is epsilon

OpenStudy (anonymous):

didnt learn this, but you couldnt find any examples in your book? you can go on calcchat if you use their book and try to find an example problem

OpenStudy (anonymous):

no it's on webassign

OpenStudy (anonymous):

the examples aren't very helpful in the book

OpenStudy (anonymous):

I haven't tried CalcChat yet, guess I will now

OpenStudy (anonymous):

yea i would find an odd number example in the book thats close to it and find the worked out solution in calcchat. good luck.

OpenStudy (anonymous):

just found one and it's the same exact problem I have

OpenStudy (anonymous):

nice, so how do you do it? or send me a link to the calcchat problem